OpenStudy (anonymous):

plzzzz need help here :D find :- (2+i3)^3

OpenStudy (anonymous):

Can you post a screenshot of your problem?

OpenStudy (anonymous):

this is my problem :D

OpenStudy (anonymous):

find \( \Huge (2+i3 )^3\)

OpenStudy (anonymous):

no i ment i , imaginary

OpenStudy (anonymous):

try expanding it out, (2+3i)(2+3i)(2+3i) keep in mind that i^2 = -1 as you go

OpenStudy (anonymous):

=8+3*4*3i+3*2*9*(-1)+(-1)*27i=-46+9i

OpenStudy (anonymous):

there is no other way ?

OpenStudy (anonymous):

lolz its kinda of easy thx guys !

OpenStudy (anonymous):

i got -46+9i

OpenStudy (anonymous):

no no i got -10-27 i

zepdrix (zepdrix):

Hmm I'm getting something different. This is how I would work on it, \[\Large\sf (2+3\mathcal i)^3\quad=\quad (2+3\mathcal i)^2(2+3\mathcal i)\]\[\Large\bf\sf =\quad (4+6\mathcal i-9)(2+3\mathcal i)\]\[\Large\sf =\quad (-5+6\mathcal i)(2+3\mathcal i)\]

OpenStudy (anonymous):

seems not right answer :O

OpenStudy (primeralph):